OpenStudy (anonymous):

log (x+12) + log x = 6

OpenStudy (anonymous):

\[\log(x(x+12))=6\]

OpenStudy (anonymous):

\[x^2+12x=10^6\] and now you are on your own because i have no idea what he answers are

OpenStudy (anonymous):

\[\color{yellow}{\text{myininaya does though}}\]

myininaya (myininaya):

its hard to read your yellow

OpenStudy (anonymous):

\[\color{blue}{\text{blue is better}}\]

OpenStudy (anonymous):

\[\color{blue}{\text{can you solve this problem?}}\]

myininaya (myininaya):

\[x^2+12x-1000000=0\] did i put in 6 zeros i cant tell thats too many zeros

myininaya (myininaya):

\[x=\frac{-12 \pm \sqrt{144-4(-1000000)}}{2}\]

OpenStudy (anonymous):

maybe it was a .... typo!

OpenStudy (anonymous):

yeah i could have done that.

myininaya (myininaya):

\[x=\frac{-12 \pm \sqrt{144+4000000}}{2}\]

myininaya (myininaya):

\[x=\frac{-12 \pm \sqrt{4000144}}{2}\]

OpenStudy (anonymous):

ok enough algebra for one night. do you know that last night some guy was trying to use linear approximation to a trig function IN DEGREES!

OpenStudy (anonymous):

question was use linear approximation to find \[\sin(44)\]!

OpenStudy (anonymous):

what kind of morons do they have teaching mathematics?

OpenStudy (anonymous):

i actually logged out, and then came back and logged in to tell the guy that he could not do it

myininaya (myininaya):

lol satellite i didn't realized we changed the subject

OpenStudy (anonymous):

oh i was just ranting. but really i was plucked

OpenStudy (anonymous):

i see a lot of idiotic questions here but that one took the cake

OpenStudy (anonymous):

then there was another problem form some idiots at "apex learning" whatever that is, that wanted the students so say the line went "down and to the left"!

myininaya (myininaya):

\[x=\frac{-12 \pm \sqrt{16}*\sqrt{250009}}{2}=\frac{-12 \pm 4 \sqrt{250009}}{2}=-6 \pm 2\sqrt{250009}\]

OpenStudy (anonymous):

ok show off. bet the question was something else. i am off to take the derivative of a trig function of degrees. i'll let you know what i get

myininaya (myininaya):

hey autuumn the only solution you have -6+2sqrt{250009{

OpenStudy (anonymous):

but i now one damn thing for sure. the derivative of sine is NOT cosine. gnight

myininaya (myininaya):

lol what?

myininaya (myininaya):

satellite have you gone mad?

OpenStudy (anonymous):

eh okay thanks everyone :)